# Beyond Prompts: Space–Time Decoupling Control-Plane Jailbreaks in LLM Structured Output

### v2

@@ 5.4. Case study: how internal alignment is broken by malicious grammar
CDA vulnerabilities are two-fold: (1) External Guardrail failure (bypassing filters) and (2) Internal Alignment failure (coercing models to generate harmful content instead of refusing). To analyze the latter, we evaluate token probability distributions using phi-3.5-moe (phi3safety2024arxiv) across three categories: refusal, safe, and jailbreak tokens.
@@
This observation strongly echoes the findings of deepalignment2025iclr, which argue that current LLM safety alignment is often “only a few tokens deep.” By leveraging grammar constraints to force these critical first few tokens—such as affirmative prefixes or malicious intent literals—CDAs effectively bypass this shallow layer of internal resistance. Once the initial safety-critical token selection is overridden via deterministic control-plane manipulation, the model’s subsequent generation behaves as if no alignment were present, highlighting a fundamental fragility in current safety-training paradigms that rely primarily on steering the start of the response.
Table 8. Summary of existing jailbreak attacks adapted from  (JBShield2025usenixsecurity), - indicates the method does not use the listed resource or lacks that capability, $\circ$ denotes white-box attack and $\bullet$ denotes black-box attack.
@@
Manually-designed IJP(jailbreak2024ccs) Human $\bullet$ - Input
@@
Template-based MasterKey(masterkey2024ndss) LLM $\bullet$ $\sim$200 Input
LLMFuzzer(llmfuzzer2024usenixsecurity) LLM $\bullet$ $\sim$500 Input
AutoDAN(autodan2024iclr) LLM $\circ$ $\sim$200 Input
PAIR(PAIR2024arxiv) LLM $\bullet$ $\sim$20 Input
TAP(TAP2024nips) LLM $\bullet$ $\sim$20 Input
StructTransform(structtransform2025arxiv) LLM $\bullet$ $\sim$3 Input
@@
Encoding-based Zulu(zulu2024arxiv) - $\bullet$ - Input
Base64(base64attack2023nips) - $\bullet$ - Input
@@ 6.1. Jailbreak Attacks on LLMs
Jailbreak attacks aim to craft malicious inputs that cause LLMs to violate their safety guidelines. Following early suggestions by Carlini et al. (align2023nips), numerous jailbreak methods have emerged. We adopt the taxonomy from JBShield (JBShield2025usenixsecurity), which categorizes attacks as manual-designed, optimization-based, template-based, linguistics-based, and encoding-based. As shown in Table 8, we extend this taxonomy with a new category, output-based attacks, which is highly relevant to our work.
Manual-designed Jailbreaks involve manually crafting malicious inputs. Notable examples include the In-the-wild Jailbreak Prompts (IJP) (jailbreak2024ccs), which document real-world attempts shared on social media.
Optimization-based Jailbreaks use automated algorithms to craft adversarial prompts. GCG (GCG2023arxiv) adds an adversarial suffix to prompts using a greedy algorithm, while SAA (SAA2025iclr) combines templates with a random search strategy. While automated, these attacks typically require white-box or logit access and a large number of queries.
Template-based Jailbreaks embed harmful requests within sophisticated, optimized templates. Methods like MasterKey (masterkey2024ndss), LLMFuzzer (llmfuzzer2024usenixsecurity), AutoDAN (autodan2024iclr), PAIR (PAIR2024arxiv), and TAP (TAP2024nips) often use other LLMs to generate or refine these templates. Concurrent work like StructTransform (structtransform2025arxiv) also targets structured generation, aligning with our findings.
Linguistics-based Jailbreaks conceal malicious intent within seemingly benign inputs using linguistic properties. For example, DrAttack (drattack2024emnlpfinding) decomposes and reconstructs prompts, while Puzzler (puzzler2024aclfinding) uses combinations of diverse clues to bypass safety mechanisms.
Encoding-based Jailbreaks obfuscate malicious prompts by translating them into less common languages (Zulu (zulu2024arxiv)) or encoding formats like Base64 (base64attack2023nips), exploiting the mismatch between the model’s high-level capabilities and its restricted safety alignment as identified by foundational work (wei2023jailbroken).
Output-based Jailbreaks is an emerging category. White-box methods like EnDec (enforcedecode2024acl) and JailMine (jailmine2024arxiv) directly manipulate model logits or token probabilities to construct harmful outputs. Closely related to our work, APT (APT2025arxiv) uses the GuidedRegex feature in structured output APIs to build a prefix tree that explicitly bans refusal tokens, forcing a harmful response.
@@ 7. Discussions
Given the significant vulnerabilities revealed by CDAs, we discuss the implications for the LLM safety landscape and propose potential mitigations. As shown in Figure 6, current auditing practices focus on two phases: input and output auditing (outauditing2024arxiv).
Input Auditing. Input-focused strategies use classifiers or small LLMs to filter prompts in parallel with generation, providing a cost-effective defense. However, CDAs easily bypass this by hiding malicious intent within the control-plane grammar while keeping the data-plane prompt benign. While auditing the grammar alongside the prompt (Dual-Plane Guard) could mitigate this, it remains impractical for complex, interleaved multi-turn attacks due to high reasoning costs and the limited capability of lightweight guard models.
Figure 9. BenignEnumAttack (red) and a benign redteaming Q-A auditing process (blue). Existing output auditing methods cannot identify their differences, so that they will be refused or returned together, causing either false positive or false negative.
Output Auditing. Output auditing filters generated content but faces steep challenges in latency, cost, and false positives (youcant2025www). Consequently, major APIs like openai and gemini often omit it. Beyond practical limits, we question if output auditing is a “silver bullet”. To test this, we developed BenignEnumAttack (Figure 9), which mimics a benign structured output (e.g., a StrongREJECT report) while embedding harmful content. This creates a semantic deadlock: since the jailbreak and a legitimate safety audit appear structurally identical, an auditor must either trigger a false positive (refusing safety research) or a false negative (leaking the jailbreak). This indicates that output-based defense has fundamental conceptual limitations beyond mere implementation costs.
Mitigation Strategies. To secure the control plane, we propose several strategies across different layers:
@@
    Lightweight Heuristics: Filters could flag suspicious Literal-to-Logic Ratios or Semantic Mismatches between prompts and schemas. However, such heuristics face similar false positive risk in Figure 9. For example, a legitimate tool for generating a technical glossary or a translation table would naturally contain many string literals and might trigger these filters.Ensuring safety without breaking legitimate utility remains a complex trade-off that requires more robust and context-aware auditing.
@@
    Grammar-Plane Auditing: Direct auditing of JSON Schemas can mitigate basic attacks like EnumAttack. However, cross-plane coordination for DictAttack remains a high-cost reasoning challenge.
@@
    Context-Aware Token Attribution: Current output auditing processes cannot distinguish between user-prefilled content and model-generated content—a vulnerability motivated by BenignEnumAttack. A more robust approach would implement token-level provenance tracking during the generation process, enabling auditing systems to differentiate between tokens originating from enum constraints (or others specified by grammars) versus those freely generated by the model. This would enable more accurate detection of attempts to manipulate the generation trajectory through constrained fields, neutralizing threats like BenignEnumAttack.
@@
    Safety-Preserving Constraints: We suggest whitelisting refusal tokens (e.g., “sorry”) to override user-defined grammars, ensuring internal safety alignment can trigger even within constrained outputs when jailbreak is absolutely detected. However, implementation challenges include balancing strict grammar adherence with safety requirements, as excessive safety tokens might disrupt valid structured outputs in certain contexts.
  5. (5)
    Integrated Safety Signaling: Co-design LLM with auditing signal tokens: we still whitelist some tokens in the constrained decoding process, but we whitelist special tokens that indicate potential auditing signals. For example, we can co-design LLMs to generate special auditing tokens (e.g., <unsafe>) when encountering harmful contexts, to enable real-time, constraint-agnostic monitoring.
@@ 8. Conclusion
In this study, we introduced the Constrained Decoding Attack (CDA), revealing a critical control-plane vulnerability in LLMs. Our evaluation across 13 models, including gpt-5 and gemini-2.5-pro, demonstrates that by weaponizing deterministic grammar constraints, attackers can bypass both internal alignment and external guardrails with near-perfect success. More importantly, our primary contribution, DictAttack, exposes a fundamental “semantic gap” in current defenses; by decoupling intent across space and time, it maintains a 75.8% ASR even against state-of-the-art jailbreak guardrails. By revealing this previously unexplored attack surface, our work contributes to developing more comprehensive security paradigms for LLMs that address safety at all stages of LLM generation.
@@
Our work identifies significant security vulnerabilities in constrained decoding mechanisms for LLMs by demonstrating how guided grammar manipulation can bypass safety guardrails. We empirically show that state-of-the-art models remain vulnerable to Constrained Decoding Attacks despite robust safety alignment efforts. The security implications are particularly concerning as CDAs exploit a fundamental gap in existing safety architectures, which primarily focus on input filtering rather than the generation process itself.
By publishing this research, we aim to enable the security community to address these vulnerabilities before they can be widely exploited. Since CDAs target low-level decoding processes exposed through structured output features, mitigating these attacks presents unique challenges distinct from traditional jailbreak defenses. We have acknowledged our findings to affected model providers like OpenAI and Gemini and propose several mitigation strategies that could reduce risk without compromising legitimate structured output functionality. We believe our analysis provides crucial insights for developing more comprehensive safety mechanisms that protect the entire generation pipeline rather than just input and output boundaries.

### v3

@@ 5.4. Case study: how internal alignment is broken by malicious grammar
Table 8. Circuit Breaker evaluation on Llama-3-8B-Instruct (Base) vs. Llama-3-8B-Instruct-$R$ (CB (Zou et al., 2024)). Top: attack ASR (%) on AdvBench[0:50] with the same gpt-4o judge as the main tables (lower is better for the defense). Bottom: benign structured-output utility (%, schema-valid rate) on JSONSchemaBench (Geng et al., 2025)’s GitHub split, 50 schemas per tier (higher preserves normal usage). Both models served locally with vllm + xgrammar (Dong et al., 2025).
Workload Base CB
Attacks (ASR)
AutoDAN-Turbo (Liu et al., 2025) (prompt-only) 82.0 6.0
EnumAttack (Ours, grammar-only) 84.0 32.0
DictAttack (Ours, grammar + dict) 24.0 14.0
[…]
Benign utility (JSONSchemaBench, GitHub split)
Trivial 84.0 84.0
Easy 76.0 74.0
Medium 84.0 84.0
Hard 48.0 54.0
DictAttack’s two-stage dictionary payload exceeds 8B-Llama’s reasoning capacity even undefended; the small CB drop is capacity-limited, not a defense effect (see §5.5).
CDA vulnerabilities are two-fold: (1) External Guardrail failure (bypassing filters) and (2) Internal Alignment failure (coercing models to generate harmful content instead of refusing). To analyze the latter, we evaluate token probability distributions using phi-3.5-moe (Haider et al., 2024) across three categories: refusal, safe, and jailbreak tokens.
@@
Table 9. Summary of existing jailbreak attacks adapted from (Zhang et al., 2025). “-” indicates the method does not use the listed resource or lacks that capability; $\circ$ denotes white-box attack, $\bullet$ denotes black-box attack.
@@
Manually-designed IJP(Shen et al., 2024) Human $\bullet$ - Input
@@
Template-based MasterKey(Deng et al., 2024) LLM $\bullet$ $\sim$200 Input
LLMFuzzer(Yu et al., 2024) LLM $\bullet$ $\sim$500 Input
AutoDAN(Liu et al., 2024a) LLM $\circ$ $\sim$200 Input
PAIR(Chao et al., 2024b) LLM $\bullet$ $\sim$20 Input
TAP(Mehrotra et al., 2024) LLM $\bullet$ $\sim$20 Input
StructTransform(Yoosuf et al., 2025) LLM $\bullet$ $\sim$3 Input
@@
Encoding-based Zulu(Yong et al., 2024) - $\bullet$ - Input
Base64(Wei et al., 2023a) - $\bullet$ - Input
@@
This observation echoes the findings of Qi et al. (2025), which argue that current LLM safety alignment is often “only a few tokens deep.” By leveraging grammar constraints to force these critical first few tokens—such as affirmative prefixes or malicious intent literals—CDAs effectively bypass this shallow layer of internal resistance. Once the initial safety-critical token selection is overridden via deterministic control-plane manipulation, the model’s subsequent generation behaves as if no alignment were present, highlighting a fundamental fragility in current safety-training paradigms that rely primarily on steering the start of the response. This view also clarifies CDAs’ relationship to prefill-style attacks (Li et al., 2025a): both exploit the model’s coherence bias once generation has been steered into a harmful, fluent trajectory. CDAs, however, go substantially deeper than prefill — the control plane can constrain not just a short prefix but a much richer, grammar-specified portion of the trajectory in a black-box manner. This raises the question of whether early-stop defenses such as Circuit Breakers (Zou et al., 2024) — which train the model to terminate (via EOS) once a harmful representation forms — can close the gap. Note that under constrained decoding the EOS token is itself masked away by the user-supplied grammar, so any early-stop signal must manifest as a degraded continuation rather than a clean stop. We answer this directly in §§ 5.5.
5.5. CDAs against Specially-Defended Models
Our main evaluation targets standard alignment-plus-guardrail stacks. Here we additionally test CDAs against Circuit Breakers (CB) (Zou et al., 2024), a representation-level defense that re-routes harmful internal states during generation. Lacking a CB-hardened flagship API, we compare the public meta-llama/Llama-3-8B-Instruct (Base) with GraySwanAI/Llama-3-8B-Instruct-RR (CB) on AdvBench[0:50], reusing the gpt-4o judge from our main tables and serving both models with vllm + xgrammar. Beyond rerunning AutoDAN-Turbo (Liu et al., 2025) as the strongest prompt-only baseline, we evaluate a combined attack — AutoDAN-Turbo’s prompt under EnumAttack’s grammar — to test whether the two attack vectors stack. To assess whether CB compromises normal usage, we measure benign utility on JSONSchemaBench (Geng et al., 2025), sampling 50 schemas from each of four difficulty tiers (trivial, easy, medium, hard) of its GitHub split.
[…]
@@ 6.1. Jailbreak Attacks on LLMs
Jailbreak attacks aim to craft malicious inputs that cause LLMs to violate their safety guidelines. Following early suggestions by Carlini et al. (Carlini et al., 2023), numerous jailbreak methods have emerged. We adopt the taxonomy from JBShield (Zhang et al., 2025), which categorizes attacks as manual-designed, optimization-based, template-based, linguistics-based, and encoding-based. As shown in Table 9, we extend this taxonomy with a new category, output-based attacks, which is highly relevant to our work.
Manual-designed Jailbreaks involve manually crafting malicious inputs. Notable examples include the In-the-wild Jailbreak Prompts (IJP) (Shen et al., 2024), which document real-world attempts shared on social media.
Optimization-based Jailbreaks use automated algorithms to craft adversarial prompts. GCG (Zou et al., 2023) adds an adversarial suffix to prompts using a greedy algorithm, while SAA (Andriushchenko et al., 2025) combines templates with a random search. While automated, these attacks require white-box or logit access and a large number of queries.
Template-based Jailbreaks embed harmful requests within sophisticated, optimized templates. Methods like MasterKey (Deng et al., 2024), LLMFuzzer (Yu et al., 2024), AutoDAN (Liu et al., 2024a), PAIR (Chao et al., 2024b), and TAP (Mehrotra et al., 2024) often use other LLMs to generate or refine these templates. Concurrent work like StructTransform (Yoosuf et al., 2025) also targets structured generation, aligning with our findings.
Linguistics-based Jailbreaks conceal malicious intent within seemingly benign inputs using linguistic properties. For example, DrAttack (Li et al., 2024a) decomposes and reconstructs prompts, while Puzzler (Chang et al., 2024) uses combinations of diverse clues to bypass alignment.
Encoding-based Jailbreaks obfuscate malicious prompts by translating them into less common languages like Zulu (Yong et al., 2024) or encoding formats like Base64 (Wei et al., 2023a), exploiting the mismatch between the model’s high-level capabilities and its restricted safety alignment as identified by foundational work (Wei et al., 2023b).
Output-based Jailbreaks form an emerging category, which we formalize as Constrained Decoding Attacks (CDA). Early white-box methods like EnDec (Zhang et al., 2024) and JailMine (Li et al., 2024b) directly manipulated model logits to enforce harmful generation. In the black-box setting, APT (Li et al., 2025b) exploits GuidedRegex to iteratively block refusal tokens via a prefix tree. However, a major limitation of these approaches is their inefficiency, often requiring iterative interactions with the target model. In contrast, our proposed attacks are highly efficient, achieving jailbreaks with $O(1)$ query complexity (a single query) to the target LLM. While DictAttack leverages LLM assistance for dictionary construction, we emphasize that this process is offline, model-agnostic, and semantically benign: it generates harmless synonyms to construct a dictionary, without requiring any adversarial feedback or access to the target model’s internal states.
@@ 7. Discussions
Given the significant vulnerabilities revealed by CDAs, we discuss the implications for the LLM safety landscape and propose potential mitigations. As shown in Figure 6, current auditing practices focus on two phases: input and output auditing (Bell and Fonseca, 2024).
Input Auditing. Input-focused strategies use classifiers or small LLMs to filter prompts in parallel with generation, providing a cost-effective defense. However, CDAs easily bypass this by hiding malicious intent within the control-plane grammar while keeping the data-plane prompt benign. While auditing the grammar alongside the prompt (Dual-Plane Guard) could mitigate this, it remains impractical for interleaved multi-turn attacks due to high auditing costs and the limited capability of lightweight guard models.
Output Auditing. Output auditing filters generated content but faces steep challenges in latency, cost, and false positives (Mai et al., 2025). Consequently, major APIs like openai and gemini often omit it. Beyond practical limits, output auditing has a conceptual ceiling against CDAs: as we illustrate with BenignEnumAttack (Figure 9), an attack response (red flow) can be made structurally indistinguishable from a legitimate StrongREJECT-style safety audit (blue flow), forcing any purely post-hoc auditor into either a false positive on legitimate safety research or a false negative on the jailbreak.
Figure 9. BenignEnumAttack (red) and a benign red-teaming Q-A auditing process (blue). Existing output auditing methods cannot identify their differences, causing either false positive or false negative.
Mitigation Strategies. We sketch four directions for securing the control plane:
@@
    Lightweight Heuristics: flag suspicious literal-to-logic ratios or prompt-schema semantic mismatches. Cheap, but inherently subject to the false-positive trade-off illustrated by Figure 9.
@@
    Grammar-Plane Auditing: directly audit user-supplied JSON Schemas. Sufficient for EnumAttack; cross-plane coordination for DictAttack remains a high-cost reasoning challenge.
@@
    Early-Stop Alignment (CB-style): train the model at the representation level to terminate or refuse harmful trajectories rather than relying on input-side filtering, as in Circuit Breakers (Zou et al., 2024). §§ 5.5 shows this materially suppresses single-vector attacks but does not close the gap once the prompt and grammar are composed (78% combined ASR), since the alignment side cannot reach the control plane.
@@
    Constraint-Aware Decoding: keep alignment-critical signals reachable under user grammars, e.g. unmaskable refusal tokens or constrained-decoding-transparent audit tokens (e.g., <unsafe>). Together with token-provenance tracking, this also neutralises BenignEnumAttack.
@@ 8. Conclusion
In this study, we introduced the Constrained Decoding Attack (CDA), revealing a critical control-plane vulnerability in LLMs. Our evaluation across 13 models, including gpt-5 and gemini-2.5-pro, demonstrates that by weaponizing deterministic grammar constraints, attackers can bypass both internal alignment and external guardrails with near-perfect success. More importantly, our primary contribution, DictAttack, exposes a fundamental “semantic gap” in current defenses: by realizing the control-to-semantic pipeline — the schema first commits malicious intent to the generation trajectory, after which the model fluently completes it — DictAttack maintains a 75.8% ASR even against SOTA jailbreak guardrails. By revealing this previously unexplored attack surface, our work contributes to developing more comprehensive security paradigms for LLMs that address safety at all stages of LLM generation.
@@
This paper exposes a previously underexplored attack surface — the LLM control plane — and shows that state-of-the-art models remain vulnerable to Constrained Decoding Attacks despite strong safety alignment. We disclosed our findings to OpenAI and Google (Gemini) in early 2025; the embargo has since passed. We additionally notified the maintainers of xgrammar, who acknowledged the vulnerability. Mitigations deployed inside closed-source GPT/Gemini stacks are not visible to us. We propose mitigation strategies that protect the full generation pipeline without disabling legitimate structured-output functionality, so that the community can close this gap before adversaries can exploit it.
Open Science
The project page here hosts self-contained proofs of concept and links to the full evaluation harness and run logs, which are available as a separate artifact under academic-only gated access; verified researchers may request access from the corresponding author. All datasets are open-access on Hugging Face. Because LLM evaluations involve sampling and model-based judging, results may vary slightly across runs.
Acknowledgments
This work is supported by the National Key Research and Development Program of China (24YFB4505603), the National Natural Science Foundation of China (62232015, 62302479, U23B2020), and the CAS Pioneer “Hundred Talents Program” (Category B, E545030000). We would like to thank anonymous reviewers for their insightful suggestions to improve this work. This paper was edited for grammar using Google Gemini (Gemini Team Google, 2023).